Amortization of a Loan

EXAMPLE EXERCISES

For the following exercises,

create an amortization table for a $21,500 car loan with 4.5% APR over 72

months.

1. What is the monthly payment?

2. How much is actually paid over the life of the loan? How much interest?

3. How much is saved if you round up to the next $50 increment? Next $100?

For the following exercises, create an amortization table for $135,000 home purchase at 4.6% fixed APR. Remember mortgages are generally paid monthly.

4. 30 years

5. 15 years

6. How do the 15-year and 30-year loans compare in total amount paid and interest paid?

7. How much is saved if you round each payment up to the nearest $100
Answer #1

Calculate the amount of installment and prepare the amortization table in the following manner:

A C D 1 2 3 4 5 6 Calculation of Amount of installment Particulars Amount Amount of loan 21500 Rate of interest per month -4.А =E23 =E24 =E25 =E26 =E27 =E28 =E29 =E30 =E31 =E32 =E33 =E34 =E35 =B24*4.5%/12 =B25*4.5%/12 =B26*4.5%/12 =B27*4.5%/12 =B28*4=E48 =E49 -E50 =E52 =E53 =E54 =E55 -E56 =E57 =E58 =E59 =E60 C =B49*4.5%/12 =B50*4.5%/12 =B51*4.5%/12 =B52*4.5%/12 =B53*4.5%/1А =E58 -E59 =E60 =E61 =E63 =E64 =E65 =E67 =E68 =E69 =B59*4.5%/12 =B60*4.5%/12 =B61*4.5%/12 =B62*4.5%/12 =B63*4.5%/12 =B64*4.5

The result will be as follows

1 2 3 4 5 6 7 А в Calculation of Amount of installment Particulars Amount Amount of loan 21500 Rate of interest per month 0.3B 18034.05223 17760.38993 |17485.70139 17209.98277 16933.2302 |16655.43982 |16376.60772 |16096.73 15815.80273 15533.82199 |15| E 8 6 42 44 45. 46 48 8 8 8 8 8 8 8 8 8 8 8 8 8 8 8 (11174.97166 141.90614371 341.29 (10875.5878 ]40.78345425 341.29 10575.65 66 в 3671.205294 3343.682314 3014.931122 2684.947114 2353.725666 2021.262137 1687.55187 1352.59019 1016.372403 678.8937993

1.

The monthly payment is $341.29

2.

Amount actually paid over the life of the loan is as follows:

1 Amount of installment 2 Number of installments 3 Total amount paid=Amount*Number 4 amount of loan Interest paid=Amount of l

3.

Amount saved after rounding to next $50 and next $100

А 1 Particulars 2 Amount of loan 3 Amount of payment 4 Rate of interest 5 Number of payments B Amount Amount 21500 21500 =350

The result will be as follows:

A 1 Particulars 2 Amount of loan 3 Amount of payment 4 Rate of interest 5 Number of payments B C Amount Amount 21500 21500 35